# Reconstruction-Free Anomaly Detection with Diffusion Models

### v2

@@ Introduction
Anomaly Score Calculation
Once we obtain the latent representation $\mathbf{z}_{T}\in\mathbb{R}^{C\times h\times w}$ for features of a given test sample $\mathbf{x}$, the most straightforward way to compute the anomaly score is via the log-likelihood under the latent Gaussian distribution. However, it often triggers the known reverse-scoring problem and fails to capture small anomalies (See supplementary materials). Given that anomalies are inherently sparse in images, we additionally propose a norm-based image-level anomaly scoring approach. Specifically, we compute the Euclidean norm over all channels $C$ for each pixel $(i,j)$ in $\mathbf{z}_{T}$, which is written as
Different from (He et al. 2024), for the image-level anomaly score, we compute the absolute difference between the maximum and minimum values of $\mathbf{z}_{T}^{\text{normed}}$. This is motivated by the observation that, even if an image is anomalous, anomalies for images are often locally and sparsely distributed within small regions, and the majority of image regions remain normal. In essence, our scheme for image-level anomaly scoring examines the difference between the highest- and the lowest-norm pixel to mitigate the influence of potential outlier pixels, thereby alleviating the reverse-scoring issue. The pixel-level anomaly score map $M_{\text{px}}$ is simply produced by interpolating $\mathbf{z}_{T}^{\text{normed}}\in\mathbb{R}^{h\times w}$ to match the desired output resolution $(H,W)$.
@@
Table 3: Quantitative results per category on MVTecAD under the single-class setting, where we report (image-level AU-ROC, pixel-level AU-ROC) and FPS. The best results per category are highlighted in bold.
Category DiffAD GLAD TransFusion DAD DDAD ScoreDD DeCoDiff InvAD (Ours)
Texture carpet (98.3, 98.1) (99.0, 98.5) (99.2, 98.9) (99.3, 99.0) (99.3, 98.7) (96.9, 98.9) (85.2, 96.3) (100, 98.0)
grid (100, 99.7) (100, 99.6) (100, 99.1) (100, 99.7) (100, 99.4) (100, 99.7) (100, 99.5) (100, 97.4)
leather (100, 99.1) (100, 99.8) (100, 99.5) (100, 99.8) (100, 99.4) (99.6, 99.3) (97.5, 98.8) (100, 98.0)
tile (100, 99.4) (100, 98.7) (99.8, 92.1) (99.9, 99.7) (100, 98.2) (98.6, 94.4) (97.4, 98.2) (100, 94.4)
wood (100, 96.7) (99.4, 98.4) (99.4, 94.5) (99.9, 96.7) (100, 95.0) (98.8, 95.1) (94.6, 97.3) (99.3, 92.3)
Object bottle (100, 98.8) (100, 98.9) (100, 97.7) (99.7, 99.1) (100, 98.7) (100, 97.9) (100, 98.7) (100, 97.7)
cable (94.6, 96.8) (99.9, 98.1) (97.9, 95.6) (99.6, 98.1) (99.4, 98.1) (96.8, 97.5) (99.2, 98.3) (99.7, 98.7)
capsule (97.5, 98.2) (99.5, 98.5) (98.5, 97.5) (98.9, 98.3) (99.4, 95.7) (96.1, 98.6) (91.0, 97.2) (99.4, 98.8)
hazelnut (100, 99.4) (100, 99.5) (100, 97.3) (100, 99.7) (100, 98.4) (99.9, 99.2) (99.5, 97.8) (99.9, 98.1)
metal nut (99.5, 99.1) (100, 98.8) (100, 96.8) (100, 99.7) (100, 99.0) (97.2, 97.9) (97.9, 97.9) (99.7, 95.4)
pill (97.7, 97.7) (98.1, 97.9) (98.3, 92.5) (99.0, 99.3) (100, 99.1) (95.3, 96.0) (96.4, 98.9) (98.7, 98.1)
screw (97.2, 99.0) (96.9, 99.1) (97.2, 99.0) (99.4, 99.0) (99.0, 99.3) (99.6, 99.6) (86.7, 98.3) (98.1, 99.6)
toothbrush (100, 99.2) (100, 99.4) (100, 98.6) (100, 98.8) (100, 98.7) (99.8, 98.3) (91.0, 97.2) (92.8, 97.8)
transistor (96.1, 93.7) (98.3, 96.2) (98.3, 93.1) (99.8, 92.9) (100, 95.3) (95.4, 95.2) (99.8, 97.9) (100, 97.9)
zipper (100, 99.0) (98.5, 97.9) (100, 97.6) (100, 99.2) (100, 98.2) (99.8, 99.3) (99.2, 99.5) (99.2, 97.4)
Avg. (98.7, 98.3) (99.3, 98.6) (99.2, 96.7) (99.7, 98.7) (99.8, 98.1) (98.2, 97.8) (95.7, 98.1) (99.1, 97.3)
FPS N/A 0.2 1.6 9.1 0.8 N/A 17 88.1
Experiments
Datasets
We use three widely-used unsupervised AD benchmarks—MVTecAD (Bergmann et al. 2019), VisA (Zou et al. 2022), and MPDD (Jezek et al. 2021)—all designed for measuring real-world AD performance on RGB industrial images. Since each benchmark provides both image- and pixel-level annotations, we can evaluate performance in terms of overall AD (image-level) as well as precise localization (pixel-level). We provide a detailed description of these datasets in the supplementary materials.
Evaluation Metrics
@@
Following (Feng et al. 2025), we employ Area Under the Receiver Operating Characteristic Curve (AU-ROC), Average Precision (AP), F1-score max (F1_max) for both image- and pixel-level AD evaluation. We also adopt the Area Under the Per-Region-Overlap (AU-PRO) (Bergmann et al. 2019) metric to examine anomaly localization performance at the pixel level. Furthermore, we apply Frames Per Second (FPS) to evaluate the efficiency of inference.
Implementations
@@
To assess the effectiveness of the proposed method, we begin by training a Diffusion Transformer (DiT)-based diffusion model (Peebles and Xie 2023) from scratch on each dataset with Equation 3, where we set $\gamma=\mathbf{1}$ as DDPM (Ho, Jain, and Abbeel 2020). To ensure fair comparisons, we employ the pre-trained EfficientNet-B4 (Tan and Le 2019) model on ImageNet-1k (Deng et al. 2009) to encode the images into feature space as OmiAD (Feng et al. 2025) and HVQ-Trans (Lu et al. 2023b). We train the diffusion model for 300 epochs across all datasets using the AdamW optimizer (Loshchilov and Hutter 2017) with a cosine learning rate schedule. The total diffusion step $T$ is set to 1000 during training using a linear noise schedule. For all experiments, we uniformly set $S=3$ to skip intermediate steps within the original $T$ steps, and create the subset $\tau_{3}=[333,666,999]$ to acclerate inversion compuatation. We conduct all throughput measurements on an NVIDIA RTX 4090 (24GB) with batch size set to 64. Note that the fluctuation in throughput measurements on different datasets is because some small test sets contain test samples less than the batch size (<64<64), where we also use a smaller batch size setting. See supplementary materials for more details.
@@
Table 4: Comparison of reconstruction-based approach (Recon.) against our inversion-based AD paradigm, under different total diffusion steps $S$ and perturbation ratios $r$, in multi-class MVTecAD with AU-ROC.
@@
Comparison with SOTA AD Methods
We compare the quantitative results against prior multi-class AD competitors, including (i) Transformer-based methods (HVQ-Trans (Lu et al. 2023b), UniAD (You et al. 2022)), (ii) reconstruction-based methods (RD4AD (Deng and Li 2022), DeSTSeg (Zhang et al. 2023c)), (iii) diffusion-based methods (DiAD (He et al. 2024), OmiAD (Feng et al. 2025)). The results are summarized in Table 2. It can be seen that our method generally yields the best image-level and highly competitive pixel-level detection performance on three datasets. Notably, our method consistently outperforms all the compared methods regarding inference efficiency by a large margin (i.e., FPS: 88.18 vs. 39.37 on MVTecAD). This can be attributed to the property of our proposed inversion-based AD paradigm, which enables high detection accuracy with fewer NFEs for acceleration. By contrast, conventional diffusion-based methods, such as DiAD (He et al. 2024), require VAE decoding for reconstruction in the pixel space, thus resulting in slower inference. It is noteworthy that our method is also faster than OmiAD whose NFE is 1, which is primarily due to the heavy diffusion backbone and its modeling strategy in high-resolution feature space.
Table 5: Generalizability evaluation of our method when incorporated into other diffusion-based AD approaches on MVTecAD and VisA with AU-ROC and FPS.
Method Image AU-ROC [%] FPS
MVTecAD VisA MVTecAD
DeCoDiff [CVPR’25] 99.3 96.1 17
DeCoDiff + InvAD 97.0 (-2.0) 93.2 (-2.8) 46 (+29)
MDM [ICML’25] 91.9 80.5 2.2
MDM + InvAD 98.2 (+6.3) 94.8 (+14.3) 63 (+60.8)
In addition, we compare our method against other prior single-class diffusion-based AD methods in Table 3, including DiffAD (Zhang et al. 2023b), GLAD (Yao et al. 2024), TransFusion (Fučka, Zavrtanik, and Skočaj 2025), DAD (Zhang et al. 2023a), DDAD (Mousakhan, Brox, and Tayyub 2023), ScoreDD (Teng et al. 2023), DeCoDiff (Beizaee et al. 2025). Similar to the multi-class scenario, our method consistently improves efficiency with competitive accuracy.
Is Inversion Effective?
@@
As our InvAD is designed for the inference stage, it can be incorporated into existing diffusion-based AD models in a plug-and-play manner. To evaluate the effectiveness, we incorporate InvAD into the inference stage for DeCoDiff (Beizaee et al. 2025) and Masked Diffusion Model (MDM) (Feng et al. 2025), and report the image-level AU-ROC and the FPS metrics on both datasets. It can be seen from Table 5 that InvAD leads to FPS gain for both methods with only a minor sacrifice in accuracy for DeCoDiff. This suggests that InvAD is model-agnostic and can be inserted as a plug-and-play module into different diffusion-based AD techniques.
Also, it is worth mentioning that for MDM, our method also leads to an evident performance improvement regarding AU-ROC. We assume that this is because, different from DeCoDiff that predicts the direction of deviation, MDM learns to optimize the noise residual during training, which suits better for our method.
@@
Table 6: Design ablation of feature space encoding modules on MVTecAD and VisA with AU-ROC and FPS.
Model Image AU-ROC [%] # Params FPS
MVTecAD VisA MVTecAD
VAE 87.1 86.9 1289M 52.6
ResNet-18 88.7 82.4 1239M 68.4
DINO 96.1 91.0 1312M 85.3
EfficientNet 99.0 96.9 1243M 88.1
@@
Table 7: Design ablation of diffusion architectures on MVTecAD and VisA with AU-ROC and FPS.
Model Image AU-ROC [%] # Params FPS
MVTecAD VisA MVTecAD
MLP 97.2 91.1 351M 241
UNet 98.0 88.4 1433M 288
DiT-base 94.4 88.4 155M 240
DiT-gigant 99.0 96.9 1223M 88
Backbone Design Ablation
@@
Our inversion-based method is considered in the pre-trained feature space rather than pixel space. In general, the selection of feature extractors heavily impacts the performance of AD methods. In Table 6, we evaluate the performance with different backbones including VAE (Rombach et al. 2022), ResNet-18 (He et al. 2016), DINO (Caron et al. 2021), and EfficientNet (Tan and Le 2019). It can be seen that stronger feature extractors (i.e., DINO-base and EfficientNet-B4) lead to better performance than those with CNN-based backbones (i.e., VAE and ResNet-18). Also, we study the influence of the design of diffusion modeling backbones on AD performance in Table 7. In particular, we compare four types of backbones: MLP, UNet, DiT-base, and DiT-gigant (Peebles and Xie 2023). In contrast to the feature encoding modules, we observe that the performance is generally similar, even with the simple MLP implementation. This suggests that our method is robust to diffusion architectures, but also relies on the quality of embeddings derived with feature extractors, as prior methods (Feng et al. 2025; Roth et al. 2021). See supplementary materials for a detailed description of backbones.
@@
Figure 2: Visualization of anomaly localization on MVTecAD (1st row), VisA (2nd row), and MPDD (3rd row). GT displays the ground-truth anomaly map.
Visualization Analysis
@@
We qualitatively evaluate the pixel-level AD performance of our inversion-based method against OmiAD (Feng et al. 2025), in Figure 2. Similar to the quantitative evaluation in Table 2, our method generally achieves comparable anomaly localization capability with OmiAD, which again, demonstrates that our method can well improve efficiency yet remain high accuracy.
Conclusion
@@
We propose a simple yet effective diffusion-based AD paradigm, “detection via noising in latent space”, to address the fundamental challenges within the conventional “detection via denoising in RGB space”. Our method, InvAD, follows a reconstruction-free manner to directly infer the final latent state with the DDIM inversion mechanism. To pursue efficiency, we simply enforce very few inversion steps to noise a test sample, yet manage to maintain high detection accuracy as the typicality can still be accurately scored. Due to the reconstruction-free design, InvAD is inherently tuning-free for the perturbation timestep. We experimentally demonstrate that our method achieves state-of-the-art detection performance, especially regarding inference efficiency.
@@
While our method can achieve satisfactory performance, it still needs more than one NFE. To further improve efficiency (i.e., only require 1 NFE), we will consider introducing diffusion distillation to compactly compress the progressive inversion process into only one single step.
Appendix A A. Dataset Details

### v3

@@ 4.1 Datasets
Table 3: Quantitative results on BMAD under the single-class setting, where we report image-level AU-ROC (det.), pixel-level AU-ROC (loc.), and FPS. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics. FPS measures the detection efficiency on the RESC dataset.
Methods BraTS2021 BTCV + LiTS RESC OCT2017 RSNA Camelyon16 mAD FPS
det. loc. det. loc. loc. loc. det. det. det.
PaDiM [ICPR’20] 79.0 94.4 50.8 90.9 75.9 91.4 91.8 77.5 67.3 79.9 20
CFlow [CVPR’20] 74.8 93.8 50.8 92.4 75.0 93.8 85.4 71.5 55.7 77.0 15
RD4AD [CVPR’22] 89.5 96.5 60.4 96.0 87.8 96.2 97.3 67.6 66.8 84.2 20
PatchCore [CVPR’22] 91.7 97.0 60.3 96.4 91.6 96.5 98.6 76.1 69.3 86.4 20
SimpleNet [CVPR’23] 82.5 94.8 72.3 97.5 76.2 77.1 94.7 69.1 62.4 80.7 10
InvAD (Ours) 90.3 97.8 60.6 97.8 93.6 96.5 99.0 82.7 67.7 87.2 88
4.2 Evaluation Metrics
@@
Following [omiad], we employ Area Under the Receiver Operating Characteristic Curve (AU-ROC), Average Precision (AP), F1-score max (F1_max) for both image- and pixel-level AD evaluation. We also adopt the Area Under the Per-Region-Overlap (AU-PRO) [MVTecAD] metric to examine pixel-level anomaly localization. The Frames Per Second (FPS) metric is further applied to study the efficiency of inference.
4.3 Implementations
@@
To assess the effectiveness of the proposed method, we begin by training a Diffusion Transformer (DiT)-based diffusion model [DiT] from scratch on each dataset with Equation 3, where we set $\gamma=\mathbf{1}$ as DDPM [DDPM]. To ensure fair comparisons, we employ the pre-trained EfficientNet-B4 [EfficientNet] model on ImageNet-1k [ImageNet] to encode the images into feature space as OmiAD [omiad] and HVQ-Trans [HVQ-Trans]. We train the diffusion model for 300 epochs across all datasets using the AdamW optimizer [AdamW] with a cosine learning rate schedule. The total diffusion step $T$ is set to 1000 during training using a linear noise schedule. For all experiments, we uniformly set $S=3$ to skip intermediate steps within the original $T$ steps, and create the subset $\tau_{3}=[333,666,999]$ to acclerate inversion compuatation. We conduct all throughput measurements on an NVIDIA RTX 4090 (24GB) with batch size set to 64. See Supp. Mat. for more details.
@@
Table 4: Comparison of reconstruction-based approach (Recon.) against our inversion-based AD paradigm, under different total diffusion steps $S$ and perturbation ratios $r$, in multi-class MVTecAD with image-level AU-ROC.
@@
4.4 Comparison with SOTA AD Methods
We compare the quantitative results against prior multi-class AD competitors, including (i) Transformer-based methods (HVQ-Trans [HVQ-Trans], UniAD [uniad]), (ii) reconstruction-based methods (RD4AD [rd4ad], DeSTSeg [destseg], SimpleNet [simplenet]), (iii) diffusion-based methods (DiAD [DiAD], OmiAD [omiad], MDM [omiad]). The results are summarized in Table 2. It can be seen that our method generally yields the best image-level and highly competitive pixel-level detection performance on three datasets. Notably, our method consistently outperforms all the compared methods regarding inference efficiency by a large margin (i.e., FPS: 88.1 v.s. 39.4 on MVTecAD). This can be attributed to the property of our proposed inversion-based AD paradigm, which enables high detection accuracy with fewer NFEs for acceleration. By contrast, conventional diffusion-based methods, such as DiAD [DiAD], require VAE decoding for reconstruction in the pixel space, thus resulting in slower inference. It is noteworthy that our method is also faster than OmiAD whose NFE is 1, which is primarily due to the heavy diffusion backbone and its modeling strategy in high-resolution feature space.
The results against diffusion-based AD models on single-class MVTecAD are displayed in Figure 1. Similarly to the multi-class setting, InvAD yields substantially reduced inference latency with state-of-the-art performance, indicating the effectiveness of inversion modeling for AD.
We next evaluate on BMAD to study the effectiveness of our InvAD on the medical domain. Here, we compare against PaDiM [padim], CFlow [cflow], RD4AD [rd4ad], PatchCore [PatchCore], and SimpleNet [simplenet]. We report image-level AU-ROC (det.), and pixel-level AU-ROC (loc.), if the ground-truth anomaly maps are provided. The results are presented in LABEL:tab:BMAD_results. It can be confirmed that InvAD achieves state-of-the-art performance across diverse BMAD medical datasets with a substantially improved inference efficiency. This further evidences the capacity and robustness of our method in handling medical AD, other than industrial inspection.
Table 5: Generalizability evaluation of our method when incorporated into other diffusion-based AD approaches on MVTecAD with image- and pixel-level AU-ROC, number of total parameters (#Params) and FPS.
Method MVTecAD FPS
det. loc.
DiAD [AAAI’24] 97.2 96.8 0.1
DiAD + InvAD 98.2 (+1.0) 97.5(+0.7) 88.1(+88)
MDM [ICML’25] 91.9 94.8 2.2
MDM + InvAD 98.2 (+6.3) 97.5 (+2.7) 63 (+60.8)
4.5 Is Inversion Effective?
@@ 4.6 Generalizability Evaluation
As our InvAD is designed for the inference stage, it can be incorporated into existing diffusion-based AD models in a plug-and-play manner. To evaluate the effectiveness, we incorporate InvAD into the inference stage for DiAD [DiAD] and MDM [omiad], and report the image- and pixel-level AU-ROC and the FPS metrics on multi-class MVTecAD. It can be seen from Table 5 that InvAD leads to an accuracy gain for both methods with substantial inference acceleration. In particular, since MDM learns to optimize the noise residual during training, it suits better for our method, as reflected in the remarkable improvement of image-AU-ROC. This suggests that InvAD is model-agnostic and can be inserted as a plug-and-play module into different diffusion-based AD techniques.
Table 6: Component ablation on multi-class MVTecAD. FDM denotes feature space diffusion models. S-Inv and M-Inv denote single- and multi-step inversion, resepectively.
Setting FDM S-Inv M-Inv mAD
A1 57.3
A2 44.9
A3 71.0
A4 (Ours) 83.7
4.7 Component Ablation
Our framework consists of two components: (i) feature space diffusion model (FDM), and (ii) multi-step inversion. To individually validate the contribution, we evaluate the AD performance of each component as well as their combinations on the multi-class MVTecAD. Table 6 reports the results of four scenarios (A1-A4). In the case of A1 (without inversion), AD performance significantly drops compared with our complete framework (A4), highlighting the importance of inversion. Furthermore, for the FDM-free case of A2 (i.e., pixel-space diffusion), single-step inversion also degrades accuracy. Combining FDM and single-step inversion (A3) substantially improves AD accuracy, and multi-step inversion (A4) brings further accuracy gains.
@@ 4.7 Component Ablation
Table 7: Design ablation of feature space encoding modules on multi-class MVTecAD with image- and pixel-level AU-ROC, number of total parameters (#Params), and FPS.
Model MVTecAD # Params FPS
det. loc.
ViT-B [ViT] 94.2 95.5 1312M 85.3
DINO-base [dino] 96.1 97.7 1312M 85.3
EfficientNet-B4 [EfficientNet] 99.0 96.9 1243M 88.1
@@
Table 8: Design ablation of diffusion architectures on multi-class MVTecAD with image- and pixel-level AU-ROC, and FPS.
Model MVTecAD # Params FPS
det. loc.
MLP [mar] 97.2 96.8 351M 241
UNet [DDPM] 98.0 97.3 1433M 288
DiT-base [DiT] 93.8 95.9 155M 240
DiT-gigant [DiT] 99.0 97.5 1223M 88
Table 9: Scoring scheme ablation under different total diffusion steps $S$ in multi-class MVTecAD with image-level AU-ROC.
Total Diffusion Steps $S$ 3 5 10 50 100 1000
NLL 96.1 95.2 93.0 89.7 89.4 89.1
Abs 98.4 98.1 95.8 83.2 82.2 82.0
NLL + Abs (Ours) 99.0 98.9 98.4 96.0 95.7 95.4
4.8 Backbone Design Ablation
@@
Our inversion-based method is considered in the pre-trained feature space rather than pixel space. In general, the selection of feature extractors heavily impacts the performance of AD methods [adpretrain, DiAD, DDAD]. In Table 7, we evaluate the performance with different backbones including ViT-B [ViT], DINO-base [dino], and EfficientNet-B4 [EfficientNet]. As expected, stronger feature extractors lead to better performance. Also, we study the influence of the design of diffusion modeling backbones on AD performance in Table 8. In particular, we compare four types of backbones: MLP, UNet, DiT-base, and DiT-gigant [DiT]. In contrast to the feature encoding modules, we observe that the performance is generally similar, even with the simple MLP implementation. This suggests that our method is robust to diffusion architectures, but also relies on the quality of embeddings derived with feature extractors, as prior methods [omiad, PatchCore]. See Supp. Mat. for a detailed description of backbones.
4.9 Scoring Scheme Ablation
As explained in Section 3.3, our anomaly scoring scheme considers both negative-log-likelihood (NLL) in the noise prior and the spatial absolute difference of Euclidean norm (Abs) to alleviate the reverse-scoring problem [DoGMKnow, OODTP]. To study the advantage of this strategy, we compare the performance in Table 9 with the conventional NLL-based scoring on MVTecAD. It can be seen that our scoring scheme (i.e., NLL + Abs) contributes to robustness to the diffusion step setting $S$, while the NLL or Abs solely do not achieve this. This demonstrates that our scoring scheme plays a key role in the proposed InvAD paradigm to realize reconstruction-free AD, where $S$ can be further set to be very small for efficiency.
@@ 4.9 Scoring Scheme Ablation
Figure 3: Visualization of anomaly localization against MDM [omiad] on MVTecAD. GT displays the ground-truth anomaly map.
4.10 Visualization Analysis
@@
We qualitatively evaluate in Figure 3 the pixel-level AD performance of InvAD against MDM [omiad]. Similar to the quantitative evaluation in Table 2, InvAD generally achieves more precise anomaly localization with fewer false positives than MDM, which again, demonstrates that our method improves efficiency and accuracy while preserving interpretability for anomaly locations without reconstruction.
5 Conclusion
@@
We propose a simple yet effective diffusion-based AD paradigm, “detection via noising in latent space”, to address the fundamental challenges within the conventional “detection via denoising in RGB space”. Our method, InvAD, is reconstruction-free and directly infers the final latent state with the DDIM inversion mechanism. To pursue efficiency, we simply enforce very few inversion steps to noise a test sample, yet manage to maintain high detection accuracy as the typicality can still be accurately scored. Due to the reconstruction-free design, InvAD is inherently tuning-free for the perturbation timestep. We experimentally demonstrate that InvAD achieves state-of-the-art detection performance, especially regarding inference efficiency.
@@
Limitation and Future Work. While our method can achieve satisfactory performance, it still needs more than one NFE. To further improve efficiency (i.e., 1 NFE), we will consider introducing diffusion distillation to compactly compress the progressive inversion process into only one single step. Also, exploring improved or task-specific inversion mechanism is an interesting future direction.
[…]
Supplementary Material
This document supplements our main paper, “InvAD: Inversion-based Reconstruction-Free Anomaly Detection with Diffusion Models”. The reference indices are consistent with those used in the main paper.
6 Dataset Details